# OpenUAS: Embeddings of Cities in Japan with Anchor Data for Cross-city Analysis of Area Usage Patterns

### v2

@@ Acknowledgements
This research was supported in part by JST CREST (JPMJCR21F2, JPMJCR22M4), NICT ( [^2] C01, 22609), NEDO (JPNP23003), CSTI SIP3(JPJ012495), and JSPS KAKENHI (22H03580, 22K18422). We would like to thank Blogwatcher Inc. for providing the valuable data used in this paper. We would like to express our gratitude to Blogwatcher, Inc.[40] for their cooperation in providing the data.
@@
To select the number of anchors and the size of data each anchor holds, we compared several variations. In this paper, we experimented with 64 to 1024 anchors and data sizes ranging from 10,000 to 80,000, comparing the $Misalignment$ for each variation. Figure 11 shows the results. The lower graph is a scatter plot with the data size of each variation on the x-axis. The results indicate that $Misalignment$ decreases as the size of the anchor data increases. However, larger anchor data sizes increase the volume of training data, leading to longer training times and greater memory usage. Considering these costs, we used anchor data of size 512x20000 in the OpenUAS dataset.
@@
We adjusted the ratio of loss backpropagation for data and anchors according to formulas 1, 2 during anchoring. In Figure 12, we examine the effects of various anchor weight functions. ‘Mixed’ simply combines the anchor data with the training data, treating the anchors in the same way as other meshes in the data without any weighting. ‘Constant’ fixes the anchor power at $\alpha=0.3$ from the beginning to the end of the training, regardless of the epoch. ‘Exponential’ decreases anchor power exponentially, as shown in the formula 2. According to the results, the ‘Exponential’ method performs better in all metrics. Additionally, ‘Mixed’ has better metrics than ‘Constant’, which may be due to the relatively large size of anchor data (512x20000) in relation to the training data in our experiments. Conversely, this suggests that if the anchor data size is small compared to the training data, simply mixing the anchor data with the training data might lead to the anchor data being ignored during learning.

### v3

@@ Acknowledgements
This research was supported in part by JST CREST (JPMJCR21F2, JPMJCR22M4), ACT-X(JPMJAX24M8), NICT ( [^2] C01, 22609), NEDO (JPNP23003), CSTI SIP3(JPJ012495), and JSPS KAKENHI (22H03580, 22K18422). We would like to express our gratitude to Blogwatcher, Inc.[41] for their cooperation in providing the data.
@@
To select the number of anchors and the size of data each anchor holds, we compared several variations. In this paper, we experimented with 64 to 1024 anchors and data sizes ranging from 10,000 to 80,000, comparing the $Misalignment$ for each variation. Figure 11 shows the results. The lower graph is a scatter plot with the data size of each variation on the x-axis. The results indicate that $Misalignment$ decreases as the size of the anchor data increases. However, larger anchor data sizes increase the volume of training data, leading to longer training times and greater memory usage. Considering these costs, we used anchor data of size $512\times 20,000$ in the OpenUAS dataset.
@@
We adjusted the ratio of loss backpropagation for data and anchors according to formulas 1, 2 during anchoring. In Figure 12, we examine the effects of various anchor weight functions. ‘Mixed’ simply combines the anchor data with the training data, treating the anchors in the same way as other areas in the data without any weighting. ‘Constant’ fixes the anchor power at $\alpha=0.3$ from the beginning to the end of the training, regardless of the epoch. ‘Exponential’ decreases anchor power exponentially, as shown in the formula 2. According to the results, the ‘Exponential’ method performs better in all metrics. Additionally, ‘Mixed’ has better metrics than ‘Constant’, which may be due to the relatively large size of anchor data ($512\times 20,000$) in compared to the training data in our experiments. Conversely, this suggests that if the anchor data size is small compared to the training data, simply mixing the anchor data with the training data might lead to the anchor data being ignored during learning.